Operations Report – November 16th

SOL: 3
Name of person filing report: Jas Purewal
Non-nominal systems: Pipe under sink leaked a bit because it isn’t attached properly. There is no seal. For now we have left it and are going to drain the sink very slowly.
Notes on non-nominal systems: NA
Spirit rover used: Yes
Hours: (before EVA) 197.7
Beginning charge: (Before EVA) 100%
Ending charge: (On return from EVA, before recharging) 62%
Currently charging: No
Opportunity rover used: No
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: No
Curiosity rover used: Yes
Hours: 209.9
Beginning charge: 100%
Ending charge: 67%
Currently charging: No
Perseverance rover used: NA
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: No
General notes on rovers: All working fine
Summary of Hab operations: put notes here
WATER USE: 15 gallons
Water (static tank): 410 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off) off
Toilet tank emptied: yes
Summary of internet: Good
Summary of suits and radios:
Summary of GreenHab operations: None
WATER USE: 5.02 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: None
Dual split: (Heat or AC, On or Off) Heat, On
Summary of RAM operations: (Tools used, work done) Two GPS units taken to main hab for use. Didn’t end up using them on today’s EVA. But we may use them tomorrow, if we don’t we will return them to the RAM. Replaced 3AAA batteries on one of the GPS units.
Summary of any observatory issues: NA
Summary of health and safety issues: Nose bleeds have improved/reduced.
Questions, concerns and requests to Mission Support: Still waiting on more trash bags please for toilet bins and for general bins. Pillows please.

Copyright © The Mars Society. All rights reserved. | Main Site